40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 9 01:36:28 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OLtrend.png
    Reply  Mon May 12 22:42:19 2014, rana, Summary, SUS, Optical Lever QPD Sum trends: they're almost all too weak OL-FB.pngarms_140512.pdf
       Reply  Mon May 12 22:52:59 2014, Jenne, Summary, SUS, Optical Lever Filters are all different OpLevFilters.pdf
       Reply  Tue May 13 13:27:58 2014, Steve, Summary, SUS, Optical Lever Servos setting table OLsums.png
          Reply  Tue May 13 17:03:05 2014, Steve, Summary, SUS, ETMX oplev laser replaced 
             Reply  Tue May 13 17:31:32 2014, Jenne, Summary, SUS, ETMX oplev laser replaced: New oplev gains set OLerr_13May2014.pdfOLfb_13May2014.pdf
                Reply  Tue May 13 22:55:57 2014, Jenne, Summary, SUS, ETMX oplev: cleanup 
             Reply  Wed May 14 10:04:06 2014, Steve, Summary, SUS, oplev laser and temperature TempRules.png
                Reply  Tue May 20 09:12:25 2014, Steve, Summary, SUS, ETMX oplev qpd board ETMXOLQPD121.jpgETMXOLQPDface121.jpgETMXolSUM.png
                   Reply  Tue May 20 15:54:39 2014, rana, Summary, SUS, ETMX oplev qpd board 
Message ID: 9975     Entry time: Tue May 20 15:54:39 2014     In reply to: 9973
Author: rana 
Type: Summary 
Category: SUS 
Subject: ETMX oplev qpd board 

 This QPD circuit (D980325-C1 ) uses the nice OP497 Quad FET opamp as the transimpedance amplifier. It has a low enough current noise, such that we can increase the resistors (R1-4) up to 100k and still be Johnson noise limited. We should also make sure that the compensation caps (C3-6) are ~2.2 nF so as to not destabilize the opamp. f_low = 1/2/pi/R/C = 730 Hz.

I will do the swap later today unless someone else gets to it first. (note: check for oscillations w/ fast scope probe after installing)


I did these modification tonight. The slideshow of some images is attached. Instead of 100k, I used 97.6k thin film, since this seemed like an oddball size that doesn't get used otherwise. I forgot to measure the dark noise of the quadrants before doing the swap, but comparing the pit/yaw/sum before/after the swap shows that the signal is basically unchanged (since pit/yaw is normalized by SUM), but that the noise is lower by a factor of a few above 100 Hz due to being above ADC noise now. Previously, it was bottoming out at ~10 prad/rHz. Since the signal is unchanged, I guess that the calibration and therefore the loop gain should not have changed either...

And the sum went up by almost 10x as expected from the resistor change.

ELOG V3.1.3-